Please help me solve this problem:
A rectangular counter is covered with 600 square tiles. The counter could have been covered with 400 tiles 1cm longer on a side. Fing the length of a side of the smaller tile.

Answer by stanbon(75887)

A rectangular counter is covered with 600 square tiles.
So the area of the counter is 600x^2 where x is the length of a side of a tile.
---------------------------------------------------------------------
The counter could have been covered with 400 tiles 1cm longer on a side.
So the area of the counter is 400(x+1)^2
-------------------------------------------------------------------------
Find the length of a side of the smaller tile.
-------------------------------------
Area 1st = Area 2nd
600x^2 = 400(x+1)^2
600x^2 = 400(x^2+2x+1)
600x^2 = 400x^2 + 800x + 400
200x^2 - 800x - 400 = 0
x^2 - 4x - 2 = 0
---------------------
x = [4 +- sqrt(16 - 4*-2)]/2
x = [4 +- sqrt(24)]/2
x = [4 +- 2sqrt(6)]/2
Positive solution:
x = 2 + 2sqrt(6) cm
=========================
Cheers,
Stan H.